40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1509     Entry time: Tue Sep 9 12:22:31 2014
Author: Evan 
Type: DailyProgress 
Category: BEAT 
Subject: Reduced south scatter shelf 

Last night, I looked at the TTFSS OUT2 on the spectrum analyzer, with the cavity unlocked and the laser PZT and broadband EOM unplugged. I believe this should give the demodulated spectrum of the RAM.

For north, the spectrum was white. For south, the spectrum was white down to 10 Hz, and below 10 Hz there was an apparent scattering shelf that rose several orders of magnitude above the white spectrum.

Tara and I went through the south path and placed an OD0.3 ND filter between various optics. We tried twisting one of the mode-matching lenses to reduce the scatter shelf, but it didn't seem to work.

This morning I went through the south path again with an OD0.5 filter, and eventually focused on the lens right after the FI. I found that I could greatly reduce the height of the scattering shelf (relative to the height of the white noise) by placing the filter downstream of lens, but not upstream of it. So I twisted this lens slightly, re-modematched into the cavity (minimal adjustment was required), and I found that the scatter shelf was reduced to less than 1 order of magnitude above the white noise.

Then I took a beat spectrum (carrier at 11.5 MHz, seems stable). The scatter shelf in the beat is not reduced by much, so I'll have to think about where to look next.

Attachment 1: noiseBudget.pdf  1.053 MB  | Hide | Hide all
noiseBudget.pdf
ELOG V3.1.3-